### v1

# “Infinitely Often” Transcendence of Gamma-Function Derivatives

Abstract.
Relatively little is known about the arithmetic properties of Gamma-function derivatives evaluated at arbitrary points $q\in\mathbb{Q}\setminus\mathbb{Z}_{\leq 0}$. In recent work, we showed that elements of the sequence $\left\{\Gamma^{\left(n\right)}\left(1\right)\right\}_{n\geq 1}$ are transcendental infinitely often. This result is now generalized to all sequences $\left\{\Gamma^{\left(n\right)}\left(q\right)\right\}_{n\geq 1}$ for $q\in\tfrac{1}{2}\mathbb{Z}\setminus\mathbb{Z}_{\leq 0}$. Furthermore, for $q\in\mathbb{Q}\setminus\tfrac{1}{2}\mathbb{Z}$ we show that at least one of the sequences $\left\{\Gamma^{\left(n\right)}\left(q\right)\right\}_{n\geq 1}$, $\left\{\Gamma^{\left(n\right)}\left(1-q\right)\right\}_{n\geq 1}$ contains infinitely many transcendental elements.
Key words and phrases:
Gamma-function derivatives; Euler-Mascheroni constant; transcendence.
1. Introduction
For any $q\in\mathbb{Q}\setminus\mathbb{Z}_{\leq 0}$, the Taylor-series expansion
$\Gamma\!\left(q+t\right)=\sum_{n=0}^{\infty}\Gamma^{\left(n\right)}\left(q\right)\frac{t^{n}}{n!}$ (1)
encodes the Gamma-function derivatives $\Gamma^{\left(n\right)}(q)$ as coefficients. Although $\Gamma\left(t\right),\>t\in\mathbb{R}\setminus\mathbb{Z}_{\leq 0}$ is well understood analytically, it appears that relatively little is known unconditionally about the arithmetic nature of $\Gamma^{\left(n\right)}(q)$ at rational points $q$.
For the base case, $q=1$, the sequence $\left\{\Gamma^{\left(n\right)}\left(1\right)\right\}_{n\geq 1}$ satisfies the recursion
$\Gamma^{\left(n\right)}\left(1\right)=-\gamma\Gamma^{\left(n-1\right)}\left(1\right)+{\displaystyle\sum_{k=1}^{n-1}}\dfrac{\left(-1\right)^{k+1}\left(n-1\right)!}{\left(n-k-1\right)!}\zeta\left(k+1\right)\Gamma^{\left(n-k-1\right)}\left(1\right),\quad n\in\mathbb{Z}_{\geq 1},$
where $\gamma=0.577215\ldots$ denotes the Euler-Mascheroni constant and $\Gamma^{\left(0\right)}\left(1\right)=1$. This system may be iterated to yield [^1] See, for example, Choi and Srivastava [1].
$\begin{array}[]{c}\Gamma^{\left(1\right)}\left(1\right)=-\gamma,\\ \Gamma^{\left(2\right)}\left(1\right)=\gamma^{2}+\zeta\left(2\right),\\ \Gamma^{\left(3\right)}\left(1\right)=-\gamma^{3}-3\zeta\left(2\right)\gamma-2\zeta\left(3\right),\\ \Gamma^{\left(4\right)}\left(1\right)=\gamma^{4}+6\zeta\left(2\right)\gamma^{2}+8\zeta\left(3\right)\gamma+\dfrac{27}{2}\zeta\left(4\right),\\ \Gamma^{\left(5\right)}\left(1\right)=-\gamma^{5}-10\zeta\left(2\right)\gamma^{3}-20\zeta\left(3\right)\gamma^{2}-\dfrac{135}{2}\zeta\left(4\right)\gamma-20\zeta\left(2\right)\zeta\left(3\right)-24\zeta\left(5\right),\\$ $\vdots\end{array}$ (2)
where: (i) each $\zeta\left(n\right),\>n\in\left\{2,4,\ldots\right\}$ can be expressed as a rational multiple of $\pi^{n}$; (ii) $\zeta\left(3\right)$ is known to be irrational (but not necessarily transcendental); and (iii) the arithmetic properties of $\gamma$ and $\zeta\left(n\right),\>n\in\left\{5,7,\ldots\right\}$ remain unsettled.
Although the polynomial forms on the right-hand side of (2) are generally thought to yield transcendental numbers, no single value $\Gamma^{\left(n\right)}\left(1\right)$ is currently known to be transcendental (or even irrational). In fact, the only unconditional transcendence results of which we are aware are:
  • at least one element of the pair $\left\{\Gamma^{\left(1\right)}\left(1\right),\Gamma^{\left(2\right)}\left(1\right)\right\}$ is transcendental (which follows immediately from the fact that $\Gamma^{\left(2\right)}\left(1\right)-\left(\Gamma^{\left(1\right)}\left(1\right)\right)^{2}=\zeta\left(2\right)=\pi^{2}/6$); and
  • the values $\Gamma^{\left(n\right)}\left(1\right)$ are transcendental infinitely often (as recently proved by Powers [2]). [^2] Fischler and Rivoal [3] provided a deep conditional transcendence result, showing that the $\Gamma^{\left(n\right)}\left(q\right)$ are transcendental for all $q\in\mathbb{Q}_{>0}$ if the unproved conjecture $\mathbf{E}\cap\mathbf{D}=\overline{\mathbb{Q}}$ is true (where $\mathbf{E}$ and $\mathbf{D}$ are rings of values taken at algebraic points by arithmetic Gevrey series of orders 1-1 and 11, respectively).
The purpose of the present work is to demonstrate that this “infinitely often” property persists (in some form) throughout the entire rational number line (excluding $\mathbb{Z}_{\leq 0}$). Our first result (Theorem 1) shows that for every integer and half-integer $q\in\tfrac{1}{2}\mathbb{Z}\setminus\mathbb{Z}_{\leq 0}$, elements of the sequence $\left\{\Gamma^{\left(n\right)}\left(q\right)\right\}_{n\geq 1}$ are transcendental infinitely often. Then, for rational $q\notin\tfrac{1}{2}\mathbb{Z}$, our second result (Theorem 2) is disjunctive, showing that at least one of the sequences $\left\{\Gamma^{\left(n\right)}\left(q\right)\right\}_{n\geq 1}$, $\left\{\Gamma^{\left(n\right)}\left(1-q\right)\right\}_{n\geq 1}$ contains infinitely many transcendental elements.
2. Results
2.1. Values of $\boldsymbol{q\in\tfrac{1}{2}\mathbb{Z}\setminus\mathbb{Z}_{\leq 0}}$
The theorem below generalizes Theorem 1(ii) of Powers [2], which addressed only the sequence $\left\{\Gamma^{\left(n\right)}\left(1\right)\right\}_{n\geq 1}$. The new result, which applies to all sequences $\left\{\Gamma^{\left(n\right)}\left(q\right)\right\}_{n\geq 1}$ for $q\in\tfrac{1}{2}\mathbb{Z}\setminus\mathbb{Z}_{\leq 0}$, is proved by analogous but more complex arguments based on classical identities for $\Gamma\left(t\right)$ along with constraints imposed by the transcendence of $\pi$.
Essentially, repeated application of the Gamma functional equation reduces $\Gamma\left(q\pm t\right)$ to $\Gamma\left(1\pm t\right)$ or $\Gamma\left(\tfrac{1}{2}\pm t\right)$, and Euler’s reflection formula gives closed forms involving $\sin(\pi t)$ or $\cos(\pi t)$, respectively. The resulting Taylor series then possess coefficients that are algebraic linear combinations of distinct powers of $\pi$, each with a non-vanishing highest-degree term. Comparing these coefficients with those obtained by products of the Taylor series in (1) – under the hypothesis that only finitely many $\Gamma^{\left(n\right)}\left(q\right)$ are transcendental – forces infinitely many distinct powers of $\pi$ to lie in a fixed finite-dimensional vector space over the algebraic numbers, thereby yielding a contradiction.
Theorem 1.
For all $q\in\tfrac{1}{2}\mathbb{Z}\setminus\mathbb{Z}_{\leq 0}$, elements of the sequence $\left\{\Gamma^{\left(n\right)}\left(q\right)\right\}_{n\geq 1}$ are transcendental infinitely often.
Proof.
Let $q\in\frac{1}{2}\mathbb{Z}\setminus\mathbb{Z}_{\leq 0}$, and choose $m\in\mathbb{Z}$ and $\xi\in\left\{1,\tfrac{1}{2}\right\}$ so that $q=\xi+m$. Repeated application of the functional equation $\Gamma\left(z+1\right)=z\,\Gamma\left(z\right)$ generates a rational function $R_{q}\left(t^{2}\right)\in\overline{\mathbb{Q}}\left(t^{2}\right)$, analytic at $t=0$ and satisfying $R_{q}\left(0\right)\in\overline{\mathbb{Q}}^{\times}$, such that
$\Gamma\left(q+t\right)\Gamma\left(q-t\right)=R_{q}\left(t^{2}\right)\Gamma\left(\xi+t\right)\Gamma\left(\xi-t\right).$ (3)
(For example, if $m\geq 0$ then we can take $R_{q}\left(t^{2}\right)=\prod_{i=0}^{m-1}\left[\left(\xi+i\right)^{2}-t^{2}\right]$, whereas if $m<0$ then we can use $R_{q}\left(t^{2}\right)=\prod_{i=m}^{-1}\left[\left(\xi+i\right)^{2}-t^{2}\right]^{-1}$. In either case, $R_{q}\left(0\right)\neq 0$ because $q\notin\mathbb{Z}_{\leq 0}$.)
Euler’s reflection formula provides the two identities,
$\Gamma\left(1+t\right)\Gamma\left(1-t\right)=\frac{\pi t}{\sin\left(\pi t\right)}$
and
$\Gamma\left(\frac{1}{2}+t\right)\Gamma\left(\frac{1}{2}-t\right)=\frac{\pi}{\cos\left(\pi t\right)},$
which together imply that
$\Gamma\left(\xi+t\right)\Gamma\left(\xi-t\right)={\displaystyle\sum_{j=0}^{\infty}}a_{j,\xi}\pi^{2\left(j+1-\xi\right)}t^{2j},$ (4)
with $a_{j,\xi}\in\mathbb{Q}^{\times}$ for all $j\geq 0$ (because the Taylor-series expansions of $\pi t/\sin\left(\pi t\right)$ and $\pi/\cos\left(\pi t\right)$ contain only non-zero rational coefficients given explicitly in terms of Bernoulli and Euler numbers, respectively). Re-expressing $R_{q}\left(t^{2}\right)$ as
$R_{q}\left(t^{2}\right)={\displaystyle\sum_{k=0}^{\infty}}b_{k,q}t^{2k},$
with $b_{k,q}\in\overline{\mathbb{Q}}$ for all $k\geq 0$ and $b_{0,q}=R_{q}\left(0\right)\neq 0$, we then see from (3) and (4) that
$\Gamma\left(q+t\right)\Gamma\left(q-t\right)={\displaystyle\sum_{j=0}^{\infty}}c_{j,q}t^{2j},$
where
$c_{j,q}={\displaystyle\sum_{k=0}^{j}}b_{k,q}a_{j-k,\xi}\pi^{2\left(j-k+1-\xi\right)}.$ (5)
In particular, for each $j\geq 0$ the coefficient $c_{j,q}$ is a $\overline{\mathbb{Q}}$-linear combination of $\left\{\pi^{2\left(1-\xi\right)},\pi^{2\left(2-\xi\right)},\ldots,\pi^{2\left(j+1-\xi\right)}\right\}$ whose highest-degree term is $b_{0,q}a_{j,\xi}\pi^{2\left(j+1-\xi\right)}$, with $b_{0,q}a_{j,\xi}\in\overline{\mathbb{Q}}^{\times}$.
Alternatively, from the Taylor-series expansions of $\Gamma\left(q+t\right)$ and $\Gamma\left(q-t\right)$ at $t=0$, we find
$\Gamma\left(q+t\right)\Gamma\left(q-t\right)=\left({\displaystyle\sum_{\ell_{1}=0}^{\infty}}\Gamma^{\left(\ell_{1}\right)}\left(q\right)\frac{t^{\ell_{1}}}{\ell_{1}!}\right)\left({\displaystyle\sum_{\ell_{2}=0}^{\infty}}\Gamma^{\left(\ell_{2}\right)}\left(q\right)\frac{\left(-t\right)^{\ell_{2}}}{\ell_{2}!}\right)$
${\displaystyle\sum_{j=0}^{\infty}}\left({\displaystyle\sum_{i=0}^{2j}}\frac{\left(-1\right)^{i}\Gamma^{\left(i\right)}\left(q\right)\Gamma^{\left(2j-i\right)}\left(q\right)}{i!\left(2j-i\right)!}\right)t^{2j},$
implying
$c_{j,q}={\displaystyle\sum_{i=0}^{2j}}\frac{\left(-1\right)^{i}\Gamma^{\left(i\right)}\left(q\right)\Gamma^{\left(2j-i\right)}\left(q\right)}{i!\left(2j-i\right)!}$ (6)
for all $j\geq 0$.
Now suppose (for purposes of contradiction) that only finitely many $\Gamma^{\left(n\right)}\left(q\right)$ are transcendental, and choose $n^{*}\in\mathbb{Z}_{\geq 1}$ such that $\Gamma^{\left(n\right)}\left(q\right)\in\overline{\mathbb{Q}}$ for all $n>n^{*}$, with
$\mathcal{V}=\textrm{span}_{\overline{\mathbb{Q}}}\left\{1,\Gamma\left(q\right),\Gamma^{\left(1\right)}\left(q\right),\Gamma^{\left(2\right)}\left(q\right),\ldots,\Gamma^{\left(n^{*}\right)}\left(q\right)\right\},$
denoting a finite-dimensional $\overline{\mathbb{Q}}$-vector space. For any $j>n^{*}$, each summand in (6) possesses at least one algebraic factor $\Gamma^{\left(i_{j}\right)}\left(q\right)$ with $i_{j}>n^{*}$ (because $i+\left(2j-i\right)=2j>2n^{*}\Longrightarrow\left(i>n^{*}\right)\vee\left(2j-i>n^{*}\right)$), and therefore lies in $\mathcal{V}$. Thus, $c_{j,q}\in\mathcal{V}$ for all $j>n^{*}$.
At the same time, it follows from (5) that the family $\left\{c_{j,q}\right\}_{j>n^{*}}$ is $\overline{\mathbb{Q}}$-linearly independent. This is because: (i) if there exists a non-trivial $\overline{\mathbb{Q}}$-linear dependence $\lambda_{1}c_{j_{1},q}+\lambda_{2}c_{j_{2},q}+\cdots+\lambda_{M}c_{j_{M},q}$ (with $j_{1}<j_{2}<\cdots<j_{M}$) for arbitrary $M\in\mathbb{Z}_{\geq 2}$, then the term $\pi^{2\left(j_{M}+1-\xi\right)}$ appears only in $c_{j_{M},q}$ (with non-zero algebraic coefficient $b_{0,q}a_{j_{M},\xi}$), forcing $\lambda_{M}=0$ (because otherwise $\pi$ would be algebraic); and (ii) iterating downward similarly forces $\lambda_{M-1}=\lambda_{M-2}=\cdots=\lambda_{1}=0$.
Consequently, for arbitrarily large $M$, the $M$ elements $\left\{c_{n^{*}+1,q},c_{n^{*}+2,q},\ldots,c_{n^{*}+M,q}\right\}$ must be $\overline{\mathbb{Q}}$-linearly independent and lie in $\mathcal{V}$, which is impossible because $\mathcal{V}$ is finite-dimensional. This contradiction confirms that $\Gamma^{\left(n\right)}\left(q\right)$ is transcendental infinitely often. ∎
$2.$ Values of $\boldsymbol{q\in\mathbb{Q}\setminus\tfrac{1}{2}\mathbb{Z}}$
The following theorem, which applies to all sequences $\left\{\Gamma^{\left(n\right)}\left(q\right)\right\}_{n\geq 1}$ for $q\in\mathbb{Q}\setminus\tfrac{1}{2}\mathbb{Z}$, is proved by arguments similar to those used in the proof of Theorem 1, but restricted by the fact that Euler’s reflection formula does not generalize to $\Gamma\left(q+t\right)\Gamma\left(q-t\right)$ in this case. For $q\in\mathbb{Q}\setminus\tfrac{1}{2}\mathbb{Z}$, the reflection formula naturally couples the points $q$ and 1q1-q, leading to a disjunctive result for the sequences $\left\{\Gamma^{\left(n\right)}\left(q\right)\right\}_{n\geq 1}$ and $\left\{\Gamma^{\left(n\right)}\left(1-q\right)\right\}_{n\geq 1}$.
Theorem 2.
For $q\in\mathbb{Q}\setminus\tfrac{1}{2}\mathbb{Z}$, at least one of the sequences $\left\{\Gamma^{\left(n\right)}\left(q\right)\right\}_{n\geq 1}$, $\left\{\Gamma^{\left(n\right)}\left(1-q\right)\right\}_{n\geq 1}$ contains infinitely many transcendental elements.
Proof.
Let $q\in\mathbb{Q}\setminus\frac{1}{2}\mathbb{Z}$, and note that Euler’s reflection formula gives
$\Gamma\left(q+t\right)\Gamma\left(1-q-t\right)=\dfrac{\pi}{\sin\left(\pi\left(q+t\right)\right)}$
${\displaystyle\sum_{j=0}^{\infty}}d_{j,q}\pi^{j+1}t^{j},$
with $d_{j,q}\in\overline{\mathbb{Q}}$ for all $j\geq 0$ (because $q\in\mathbb{Q}$ implies $\sin\left(\pi q\right),\cos\left(\pi q\right)\in\overline{\mathbb{Q}}$ and each $t$-derivative introduces exactly one factor of $\pi$ by the chain rule) and $d_{j,q}\neq 0$ for infinitely many $j$ (because $\pi/\sin\left(\pi\left(q+t\right)\right)$ is not a polynomial in $t$). Moreover, from the Taylor-series expansions of $\Gamma\left(q+t\right)$ and $\Gamma\left(1-q-t\right)$ at $t=0$, we see
$\Gamma\left(q+t\right)\Gamma\left(1-q-t\right)=\left({\displaystyle\sum_{\ell_{1}=0}^{\infty}}\Gamma^{\left(\ell_{1}\right)}\left(q\right)\frac{t^{\ell_{1}}}{\ell_{1}!}\right)\left({\displaystyle\sum_{\ell_{2}=0}^{\infty}}\Gamma^{\left(\ell_{2}\right)}\left(1-q\right)\frac{\left(-t\right)^{\ell_{2}}}{\ell_{2}!}\right)$
${\displaystyle\sum_{j=0}^{\infty}}\left({\displaystyle\sum_{i=0}^{j}}\frac{\left(-1\right)^{j-i}\Gamma^{\left(i\right)}\left(q\right)\Gamma^{\left(j-i\right)}\left(1-q\right)}{i!\left(j-i\right)!}\right)t^{j},$
implying
$d_{j,q}\pi^{j+1}={\displaystyle\sum_{i=0}^{j}}\frac{\left(-1\right)^{j-i}\Gamma^{\left(i\right)}\left(q\right)\Gamma^{\left(j-i\right)}\left(1-q\right)}{i!\left(j-i\right)!}$ (7)
for all $j\geq 0$.
Now suppose (for purposes of contradiction) that both sequences $\left\{\Gamma^{\left(n\right)}\left(q\right)\right\}_{n\geq 1}$ and $\left\{\Gamma^{\left(n\right)}\left(1-q\right)\right\}_{n\geq 1}$ contain only finitely many transcendental elements, and choose $n^{*}\in\mathbb{Z}_{\geq 1}$ such that both $\Gamma^{\left(n\right)}\left(q\right)\in\overline{\mathbb{Q}}$ and $\Gamma^{\left(n\right)}\left(1-q\right)\in\overline{\mathbb{Q}}$ for all $n>n^{*}$, with
$\mathcal{W}=\textrm{span}_{\overline{\mathbb{Q}}}\left\{1,\Gamma\left(q\right),\Gamma\left(1-q\right),\Gamma^{\left(1\right)}\left(q\right),\Gamma^{\left(1\right)}\left(1-q\right),\Gamma^{\left(2\right)}\left(q\right),\Gamma^{\left(2\right)}\left(1-q\right),\ldots,\Gamma^{\left(n^{*}\right)}\left(q\right),\Gamma^{\left(n^{*}\right)}\left(1-q\right)\right\}.$
For any $j>2n^{*}$, each summand in (7) has at least one algebraic factor (because $i+\left(j-i\right)=j>2n^{*}$), and therefore lies in $\mathcal{W}$. Thus, $d_{j,q}\pi^{j+1}\in\mathcal{W}$ for all $j>2n^{*}$. However, since $d_{j,q}\neq 0$ for infinitely many $j$, it follows that $\mathcal{W}$ contains infinitely many distinct powers of $\pi$, forcing a non-trivial polynomial relation over $\overline{\mathbb{Q}}$ that is satisfied by $\pi$. This contradicts the transcendence of $\pi$, thereby confirming that at least one of the sequences $\left\{\Gamma^{\left(n\right)}\left(q\right)\right\}_{n\geq 1}$, $\left\{\Gamma^{\left(n\right)}\left(1-q\right)\right\}_{n\geq 1}$ contains infinitely many transcendental elements. ∎
3. Conclusion
The present study has pursued a focused objective: generalizing the “infinitely often” transcendence results of Powers [2] to $\Gamma^{\left(n\right)}\left(q\right)$ for values $q\in\mathbb{Q}\setminus\mathbb{Z}_{\leq 0}$ beyond $q=1$. This was accomplished for $q\in\tfrac{1}{2}\mathbb{Z}\setminus\mathbb{Z}_{\leq 0}$ in Theorem 1, and partially accomplished for $q\in\mathbb{Q}\setminus\tfrac{1}{2}\mathbb{Z}$ in Theorem 2.
Natural directions for further research include: (i) improving Theorem 2 by removing the disjunctive limitation; and (ii) ascertaining lower bounds on the asymptotic densities of the transcendental terms $\Gamma^{\left(n\right)}\left(q\right)$ for $q\in\tfrac{1}{2}\mathbb{Z}\setminus\mathbb{Z}_{\leq 0}$. The latter objective, although challenging, appears feasible because Theorem 1(iv) of Powers [2] was able to provide such a lower bound for the sequence $\left\{\Gamma^{\left(n\right)}\left(1\right)\right\}_{n\geq 1}$, conditional on a conjecture rigorously certified by computer for all $n\in\left\{2,3,\ldots,26\right\}$.

### v2

@@
 Relatively little is known about the arithmetic properties of Gamma-function derivatives evaluated at arbitrary points $q\in\mathbb{Q}\setminus\mathbb{Z}_{\leq 0}$.
-In recent work, we showed that elements of the sequence $\left\{\Gamma^{\left(n\right)}\left(1\right)\right\}_{n\geq 1}$ are transcendental infinitely often.
-This result is now generalized to all sequences $\left\{\Gamma^{\left(n\right)}\left(q\right)\right\}_{n\geq 1}$ for $q\in\tfrac{1}{2}\mathbb{Z}\setminus\mathbb{Z}_{\leq 0}$.
-Furthermore, for $q\in\mathbb{Q}\setminus\tfrac{1}{2}\mathbb{Z}$ we show that at least one of the sequences $\left\{\Gamma^{\left(n\right)}\left(q\right)\right\}_{n\geq 1}$, $\left\{\Gamma^{\left(n\right)}\left(1-q\right)\right\}_{n\geq 1}$ contains infinitely many transcendental elements.
+In recent work, we showed that the sequence $\left\{\Gamma^{\left(n\right)}\left(1\right)\right\}_{n\geq 1}$ contains transcendental elements infinitely often.
+That result is now generalized to all sequences $\left\{\Gamma^{\left(n\right)}\left(q\right)\right\}_{n\geq 1}$ for $q\in\tfrac{1}{2}\mathbb{Z}\setminus\mathbb{Z}_{\leq 0}$.
+Moreover, for all such $q$ we derive a lower bound, $\beta\left(N\right)=\max\left\{0,\sqrt{N}-5/2\right\}/N$, for the density of transcendental elements $\Gamma^{\left(n\right)}\left(q\right)$ among $n\in\left\{1,2,\ldots,N\right\}$.
+For $q\in\mathbb{Q}\setminus\tfrac{1}{2}\mathbb{Z}$, we find the somewhat weaker result that at least one of the sequences $\left\{\Gamma^{\left(n\right)}\left(q\right)\right\}_{n\geq 1}$, $\left\{\Gamma^{\left(n\right)}\left(1-q\right)\right\}_{n\geq 1}$ contains infinitely many transcendental elements.
 Key words and phrases:
-Gamma-function derivatives; Euler-Mascheroni constant; transcendence.
+Gamma-function derivatives; Euler-Mascheroni constant; transcendence; density bounds.
 1. Introduction
@@
 $\Gamma\left(q+t\right)=\sum_{n=0}^{\infty}\Gamma^{\left(n\right)}\left(q\right)\frac{t^{n}}{n!}$ (1)
-encodes the Gamma-function derivatives $\Gamma^{\left(n\right)}(q)$ as coefficients.
-Although $\Gamma\left(t\right),\>t\in\mathbb{R}\setminus\mathbb{Z}_{\leq 0}$ is well understood analytically, it appears that relatively little is known unconditionally about the arithmetic nature of $\Gamma^{\left(n\right)}(q)$ at rational points $q$.
-For the base case, $q=1$, the sequence $\left\{\Gamma^{\left(n\right)}\left(1\right)\right\}_{n\geq 1}$ satisfies the recursion
-$\Gamma^{\left(n\right)}\left(1\right)=-\gamma\Gamma^{\left(n-1\right)}\left(1\right)+{\displaystyle\sum_{k=1}^{n-1}}\dfrac{\left(-1\right)^{k+1}\left(n-1\right)!}{\left(n-k-1\right)!}\zeta\left(k+1\right)\Gamma^{\left(n-k-1\right)}\left(1\right),\quad n\in\mathbb{Z}_{\geq 1},$
-where $\gamma=0.577215\ldots$ denotes the Euler-Mascheroni constant and $\Gamma^{\left(0\right)}\left(1\right)=1$. This system may be iterated to yield [^1] See, for example, Choi and Srivastava [1].
+(for sufficiently small $\left|t\right|$) encodes the Gamma-function derivatives $\Gamma^{\left(n\right)}(q)$ as coefficients.
+Although $\Gamma\left(z\right)$ is well understood analytically for $z\in\mathbb{R}\setminus\mathbb{Z}_{\leq 0}$ (see, e.g., Artin [1]), it appears that relatively little is known unconditionally about the arithmetic nature of $\Gamma^{\left(n\right)}\left(q\right)$ at rational points $q$ (see Rivoal [2, 3]).
+For the cases of $q\in\left\{1,\tfrac{1}{2}\right\}$, the sequence $\left\{\Gamma^{\left(n\right)}\left(q\right)\right\}_{n\geq 1}$ satisfies the recursions
+$\Gamma^{\left(n\right)}\left(1\right)=-\gamma\Gamma^{\left(n-1\right)}\left(1\right)+{\displaystyle\sum_{k=1}^{n-1}}\dfrac{\left(-1\right)^{k+1}\left(n-1\right)!}{\left(n-k-1\right)!}\zeta\left(k+1\right)\Gamma^{\left(n-k-1\right)}\left(1\right),$
+and
+$\Gamma^{\left(n\right)}\left(\tfrac{1}{2}\right)=-g\Gamma^{\left(n-1\right)}\left(\tfrac{1}{2}\right)+{\displaystyle\sum_{k=1}^{n-1}}\dfrac{\left(-1\right)^{k+1}\left(n-1\right)!}{\left(n-k-1\right)!}\left(2^{k+1}-1\right)\zeta\left(k+1\right)\Gamma^{\left(n-k-1\right)}\left(\tfrac{1}{2}\right),$
+respectively, where $\gamma=0.577215\ldots$ denotes the Euler-Mascheroni constant, $g=\gamma+2\ln\left(2\right)$, and $\Gamma^{\left(0\right)}\left(q\right)\equiv\Gamma\left(q\right)$.
+These identities can be iterated to yield the systems [^1] See, e.g., Choi and Srivastava [4].
 $\begin{array}[]{c}\Gamma^{\left(1\right)}\left(1\right)=-\gamma,\\ \Gamma^{\left(2\right)}\left(1\right)=\gamma^{2}+\zeta\left(2\right),\\ \Gamma^{\left(3\right)}\left(1\right)=-\gamma^{3}-3\zeta\left(2\right)\gamma-2\zeta\left(3\right),\\ \Gamma^{\left(4\right)}\left(1\right)=\gamma^{4}+6\zeta\left(2\right)\gamma^{2}+8\zeta\left(3\right)\gamma+\dfrac{27}{2}\zeta\left(4\right),\\ \Gamma^{\left(5\right)}\left(1\right)=-\gamma^{5}-10\zeta\left(2\right)\gamma^{3}-20\zeta\left(3\right)\gamma^{2}-\dfrac{135}{2}\zeta\left(4\right)\gamma-20\zeta\left(2\right)\zeta\left(3\right)-24\zeta\left(5\right),\\$ $\vdots\end{array}$ (2)
-where: (i) each $\zeta\left(n\right),\>n\in\left\{2,4,\ldots\right\}$ can be expressed as a rational multiple of $\pi^{n}$; (ii) $\zeta\left(3\right)$ is known to be irrational (but not necessarily transcendental); and (iii) the arithmetic properties of $\gamma$ and $\zeta\left(n\right),\>n\in\left\{5,7,\ldots\right\}$ remain unsettled.
-Although the polynomial forms on the right-hand side of (2) are generally thought to yield transcendental numbers, no single value $\Gamma^{\left(n\right)}\left(1\right)$ is currently known to be transcendental (or even irrational).
-In fact, the only unconditional transcendence results of which we are aware are:
-  • at least one element of the pair $\left\{\Gamma^{\left(1\right)}\left(1\right),\Gamma^{\left(2\right)}\left(1\right)\right\}$ is transcendental (which follows immediately from the fact that $\Gamma^{\left(2\right)}\left(1\right)-\left(\Gamma^{\left(1\right)}\left(1\right)\right)^{2}=\zeta\left(2\right)=\pi^{2}/6$); and
-  • the values $\Gamma^{\left(n\right)}\left(1\right)$ are transcendental infinitely often (as recently proved by Powers [2]).
-[^2] Fischler and Rivoal [3] provided a deep conditional transcendence result, showing that the $\Gamma^{\left(n\right)}\left(q\right)$ are transcendental for all $q\in\mathbb{Q}_{>0}$ if the unproved conjecture $\mathbf{E}\cap\mathbf{D}=\overline{\mathbb{Q}}$ is true (where $\mathbf{E}$ and $\mathbf{D}$ are rings of values taken at algebraic points by arithmetic Gevrey series of orders 1-1 and 11, respectively).
-The purpose of the present work is to demonstrate that this “infinitely often” property persists (in some form) throughout the entire rational number line (excluding $\mathbb{Z}_{\leq 0}$).
-Our first result (Theorem 1) shows that for every integer and half-integer $q\in\tfrac{1}{2}\mathbb{Z}\setminus\mathbb{Z}_{\leq 0}$, elements of the sequence $\left\{\Gamma^{\left(n\right)}\left(q\right)\right\}_{n\geq 1}$ are transcendental infinitely often.
-Then, for rational $q\notin\tfrac{1}{2}\mathbb{Z}$, our second result (Theorem 2) is disjunctive, showing that at least one of the sequences $\left\{\Gamma^{\left(n\right)}\left(q\right)\right\}_{n\geq 1}$, $\left\{\Gamma^{\left(n\right)}\left(1-q\right)\right\}_{n\geq 1}$ contains infinitely many transcendental elements.
-2. Results
-2.1. Values of $\boldsymbol{q\in\tfrac{1}{2}\mathbb{Z}\setminus\mathbb{Z}_{\leq 0}}$
-The theorem below generalizes Theorem 1(ii) of Powers [2], which addressed only the sequence $\left\{\Gamma^{\left(n\right)}\left(1\right)\right\}_{n\geq 1}$.
-The new result, which applies to all sequences $\left\{\Gamma^{\left(n\right)}\left(q\right)\right\}_{n\geq 1}$ for $q\in\tfrac{1}{2}\mathbb{Z}\setminus\mathbb{Z}_{\leq 0}$, is proved by analogous but more complex arguments based on classical identities for $\Gamma\left(t\right)$ along with constraints imposed by the transcendence of $\pi$.
-Essentially, repeated application of the Gamma functional equation reduces $\Gamma\left(q\pm t\right)$ to $\Gamma\left(1\pm t\right)$ or $\Gamma\left(\tfrac{1}{2}\pm t\right)$, and Euler’s reflection formula gives closed forms involving $\sin(\pi t)$ or $\cos(\pi t)$, respectively.
+and
+$\begin{array}[]{c}\Gamma^{\left(1\right)}\left(\tfrac{1}{2}\right)=-\sqrt{\pi}g,\\ \Gamma^{\left(2\right)}\left(\tfrac{1}{2}\right)=\sqrt{\pi}\left(g^{2}+3\zeta\left(2\right)\right),\\ \Gamma^{\left(3\right)}\left(\tfrac{1}{2}\right)=-\sqrt{\pi}\left(g^{3}+9\zeta\left(2\right)g+14\zeta\left(3\right)\right),\\ \Gamma^{\left(4\right)}\left(\tfrac{1}{2}\right)=\sqrt{\pi}\left(g^{4}+18\zeta\left(2\right)g^{2}+56\zeta\left(3\right)g+\dfrac{315}{2}\zeta\left(4\right)\right),\\ \Gamma^{\left(5\right)}\left(\tfrac{1}{2}\right)=-\sqrt{\pi}\left(g^{5}+30\zeta\left(2\right)g^{3}+140\zeta\left(3\right)g^{2}+\dfrac{1575}{2}\zeta\left(4\right)g+420\zeta\left(2\right)\zeta\left(3\right)+744\zeta\left(5\right)\right),\\$ $\vdots\end{array}$ (3)
+where: (i) each $\zeta\left(n\right),\>n\in\left\{2,4,\ldots\right\}$ is a rational multiple of $\pi^{n}$; (ii) $\zeta\left(3\right)$ is known to be irrational (but not necessarily transcendental); [^2] See Apéry [5] and Beukers [6].
+and (iii) the arithmetic properties of $\gamma$ and $\zeta\left(n\right),\>n\in\left\{5,7,\ldots\right\}$ remain unsettled.
+[^3] See Lagarias [7], Rivoal [2, 3], Ball and Rivoal [8], and Fischler, Sprang, and Zudilin [9].
+Although the polynomial forms on the right-hand sides of (2) and (3) are generally thought to yield transcendental numbers, no specific value of $\Gamma^{\left(n\right)}\left(1\right)$ or $\Gamma^{\left(n\right)}\left(\tfrac{1}{2}\right)$, for $n\in\mathbb{Z}_{\geq 1}$, is currently known to be transcendental (or even irrational).
+In fact, we are aware of only two unconditional transcendence results for these systems:
+[…]
+[…]
+In the present work, we investigate the “infinitely often” transcendence of $\Gamma^{\left(n\right)}\left(q\right)$ for $q\in\mathbb{Q}\setminus\mathbb{Z}_{\leq 0}$.
+First, Theorem 1 demonstrates that for every positive integer and every half-integer (i.e., for all $q\in\tfrac{1}{2}\mathbb{Z}\setminus\mathbb{Z}_{\leq 0}$), the sequence $\left\{\Gamma^{\left(n\right)}\left(q\right)\right\}_{n\geq 1}$ contains infinitely many transcendental elements.
+This analysis leads directly to Theorem 2, which shows that for all such $q$ the density of transcendental elements $\Gamma^{\left(n\right)}\left(q\right)$ among $n\in\left\{1,2,\ldots,N\right\}$ is bounded below by
+$\beta\left(N\right)=\dfrac{\max\left\{0,\sqrt{N}-5/2\right\}}{N},$
+where $\beta\left(N\right)\asymp N^{-1/2}\rightarrow 0$ as $N\rightarrow\infty$.
+Finally, for $q\in\mathbb{Q}\setminus\tfrac{1}{2}\mathbb{Z}$, Theorem 3 provides a weaker (disjunctive) counterpart to Theorem 1, asserting that at least one of the sequences $\left\{\Gamma^{\left(n\right)}\left(q\right)\right\}_{n\geq 1}$, $\left\{\Gamma^{\left(n\right)}\left(1-q\right)\right\}_{n\geq 1}$ contains infinitely many transcendental elements.
+2. Values of $q\in\tfrac{1}{2}\mathbb{Z}\setminus\mathbb{Z}_{\leq 0}$
+2.1. “Infinitely Often” Transcendental Derivatives
+We begin by showing that for all $q\in\tfrac{1}{2}\mathbb{Z}\setminus\mathbb{Z}_{\leq 0}$, transcendental elements appear infinitely often in the sequence $\left\{\Gamma^{\left(n\right)}\left(q\right)\right\}_{n\geq 1}$.
+This is proved by arguments based on classical identities for $\Gamma\left(t\right)$ along with constraints imposed by the transcendence of $\pi$.
+Essentially, repeated application of the Gamma functional equation reduces $\Gamma\left(q\pm t\right)$ to either $\Gamma\left(1\pm t\right)$ or $\Gamma\left(\tfrac{1}{2}\pm t\right)$, and Euler’s reflection formula, in conjunction with the functional equation, gives closed forms involving $\sin(\pi t)$ or $\cos(\pi t)$, respectively.
 The resulting Taylor series then possess coefficients that are algebraic linear combinations of distinct powers of $\pi$, each with a non-vanishing highest-degree term.
@@
 In either case, $R_{q}\left(0\right)\neq 0$ because $q\notin\mathbb{Z}_{\leq 0}$.)
-Euler’s reflection formula provides the two identities,
+Now note that Euler’s reflection formula, in conjunction with the functional equation, provides the identities
 $\Gamma\left(1+t\right)\Gamma\left(1-t\right)=\frac{\pi t}{\sin\left(\pi t\right)}$
 and
-$\Gamma\left(\frac{1}{2}+t\right)\Gamma\left(\frac{1}{2}-t\right)=\frac{\pi}{\cos\left(\pi t\right)},$
+$\Gamma\left(\tfrac{1}{2}+t\right)\Gamma\left(\tfrac{1}{2}-t\right)=\frac{\pi}{\cos\left(\pi t\right)},$
 which together imply that
 $\Gamma\left(\xi+t\right)\Gamma\left(\xi-t\right)={\displaystyle\sum_{j=0}^{\infty}}a_{j,\xi}\pi^{2\left(j+1-\xi\right)}t^{2j},$ (5)
-with $a_{j,\xi}\in\mathbb{Q}^{\times}$ for all $j\geq 0$ (because the Taylor-series expansions of $\pi t/\sin\left(\pi t\right)$ and $\pi/\cos\left(\pi t\right)$ contain only non-zero rational coefficients given explicitly in terms of Bernoulli and Euler numbers, respectively).
+with $a_{j,\xi}\in\mathbb{Q}^{\times}$ for all $j\geq 0$ (because the Taylor-series expansions of $\pi t/\sin\left(\pi t\right)$ and $\pi/\cos\left(\pi t\right)$ contain only even powers with non-zero rational coefficients – given explicitly in terms of Bernoulli and Euler numbers, respectively).
 Re-expressing $R_{q}\left(t^{2}\right)$ as
-$R_{q}\left(t^{2}\right)={\displaystyle\sum_{k=0}^{\infty}}b_{k,q}t^{2k},$
-with $b_{k,q}\in\overline{\mathbb{Q}}$ for all $k\geq 0$ and $b_{0,q}=R_{q}\left(0\right)\neq 0$, we then see from (3) and (4) that
+$R_{q}\left(t^{2}\right)={\displaystyle\sum_{k=0}^{\infty}}b_{k,q}t^{2k},$ (6)
+with $b_{k,q}\in\overline{\mathbb{Q}}$ for all $k\geq 0$ and $b_{0,q}=R_{q}\left(0\right)\neq 0$, we then see from (4), (5), and (6) that
 $\Gamma\left(q+t\right)\Gamma\left(q-t\right)={\displaystyle\sum_{j=0}^{\infty}}c_{j,q}t^{2j},$
@@
 denoting a finite-dimensional $\overline{\mathbb{Q}}$-vector space.
-For any $j>n^{*}$, each summand in (6) possesses at least one algebraic factor $\Gamma^{\left(i_{j}\right)}\left(q\right)$ with $i_{j}>n^{*}$ (because $i+\left(2j-i\right)=2j>2n^{*}\Longrightarrow\left(i>n^{*}\right)\vee\left(2j-i>n^{*}\right)$), and therefore lies in $\mathcal{V}$.
+For any $j>n^{*}$, each summand in (8) possesses at least one algebraic factor $\Gamma^{\left(i_{j}\right)}\left(q\right)$ for some $i_{j}>n^{*}$ (because $i+\left(2j-i\right)=2j>2n^{*}\Longrightarrow\left(i>n^{*}\right)\vee\left(2j-i>n^{*}\right)$), and therefore lies in $\mathcal{V}$.
 Thus, $c_{j,q}\in\mathcal{V}$ for all $j>n^{*}$.
 At the same time, it follows from (7) that the family $\left\{c_{j,q}\right\}_{j>n^{*}}$ is $\overline{\mathbb{Q}}$-linearly independent.
-This is because: (i) if there exists a non-trivial $\overline{\mathbb{Q}}$-linear dependence $\lambda_{1}c_{j_{1},q}+\lambda_{2}c_{j_{2},q}+\cdots+\lambda_{M}c_{j_{M},q}$ (with $j_{1}<j_{2}<\cdots<j_{M}$) for arbitrary $M\in\mathbb{Z}_{\geq 2}$, then the term $\pi^{2\left(j_{M}+1-\xi\right)}$ appears only in $c_{j_{M},q}$ (with non-zero algebraic coefficient $b_{0,q}a_{j_{M},\xi}$), forcing $\lambda_{M}=0$ (because otherwise $\pi$ would be algebraic); and (ii) iterating downward similarly forces $\lambda_{M-1}=\lambda_{M-2}=\cdots=\lambda_{1}=0$.
-Consequently, for arbitrarily large $M$, the $M$ elements $\left\{c_{n^{*}+1,q},c_{n^{*}+2,q},\ldots,c_{n^{*}+M,q}\right\}$ must be $\overline{\mathbb{Q}}$-linearly independent and lie in $\mathcal{V}$, which is impossible because $\mathcal{V}$ is finite-dimensional.
+This is because: (i) if there exists a non-trivial $\overline{\mathbb{Q}}$-linear dependence $\lambda_{1}c_{j_{1},q}+\lambda_{2}c_{j_{2},q}+\cdots+\lambda_{\mu}c_{j_{\mu},q}=0$ (with $j_{1}<j_{2}<\cdots<j_{\mu}$) for arbitrary $\mu\in\mathbb{Z}_{\geq 2}$, then the term $\pi^{2\left(j_{\mu}+1-\xi\right)}$ appears only in $c_{j_{\mu},q}$ (with non-zero algebraic coefficient $b_{0,q}a_{j_{\mu},\xi}$), forcing $\lambda_{\mu}=0$ (because otherwise $\pi$ would be algebraic); and (ii) iterating downward similarly forces $\lambda_{\mu-1}=\lambda_{\mu-2}=\cdots=\lambda_{1}=0$.
+Consequently, for arbitrarily large $\mu$, the $\mu$ elements $\left\{c_{n^{*}+1,q},c_{n^{*}+2,q},\ldots,c_{n^{*}+\mu,q}\right\}$ must be $\overline{\mathbb{Q}}$-linearly independent and lie in $\mathcal{V}$, which is impossible because $\mathcal{V}$ is finite-dimensional.
 This contradiction confirms that $\Gamma^{\left(n\right)}\left(q\right)$ is transcendental infinitely often.
 ∎
-$2.$ Values of $\boldsymbol{q\in\mathbb{Q}\setminus\tfrac{1}{2}\mathbb{Z}}$
-The following theorem, which applies to all sequences $\left\{\Gamma^{\left(n\right)}\left(q\right)\right\}_{n\geq 1}$ for $q\in\mathbb{Q}\setminus\tfrac{1}{2}\mathbb{Z}$, is proved by arguments similar to those used in the proof of Theorem 1, but restricted by the fact that Euler’s reflection formula does not generalize to $\Gamma\left(q+t\right)\Gamma\left(q-t\right)$ in this case.
-For $q\in\mathbb{Q}\setminus\tfrac{1}{2}\mathbb{Z}$, the reflection formula naturally couples the points $q$ and 1q1-q, leading to a disjunctive result for the sequences $\left\{\Gamma^{\left(n\right)}\left(q\right)\right\}_{n\geq 1}$ and $\left\{\Gamma^{\left(n\right)}\left(1-q\right)\right\}_{n\geq 1}$.
+2.2. Lower Bound for Transcendental Densities
+Let
+$\delta_{q}\left(N\right)=\dfrac{\#\left\{1\leq n\leq N:\Gamma^{\left(n\right)}\left(q\right)\textrm{ is transcendental}\right\}}{N},\quad N\in\mathbb{Z}_{\geq 1}$
+denote the density of transcendental elements $\Gamma^{\left(n\right)}\left(q\right)$ among $n\in\left\{1,2,\ldots,N\right\}$, with $q\in\frac{1}{2}\mathbb{Z}\setminus\mathbb{Z}_{\leq 0}$ fixed.
+We now provide a lower bound for this density.
+Theorem 2.
+For any $q\in\frac{1}{2}\mathbb{Z}\setminus\mathbb{Z}_{\leq 0}$,
+$\delta_{q}\left(N\right)\geq\dfrac{\max\left\{0,\left\lceil\sqrt{2\left\lfloor N/2\right\rfloor+9/4}-5/2\right\rceil\right\}}{N}\geq\dfrac{\max\left\{0,\sqrt{N}-5/2\right\}}{N}=\beta\left(N\right),$
+where $\beta\left(N\right)\asymp N^{-1/2}$ as $N\rightarrow\infty$.
+Proof.
+Fixing $q\in\frac{1}{2}\mathbb{Z}\setminus\mathbb{Z}_{\leq 0}$ and $N\in\mathbb{Z}_{\geq 1}$, consider the set of coefficients $\left\{c_{j,q}\right\}$ for $j\in\left\{0,1,\ldots,\left\lfloor N/2\right\rfloor\right\}$.
+From (8), we know that each $c_{j,q}$ is a $\overline{\mathbb{Q}}$-linear combination of products $\Gamma^{\left(u\right)}\left(q\right)\Gamma^{\left(v\right)}\left(q\right)$, with $0\leq u,v\leq N$.
+Assume that the number of transcendental elements $\Gamma^{\left(n\right)}\left(q\right)$ among $n\in\left\{1,2,\ldots,N\right\}$ is given by $N\delta_{q}\left(N\right)=\tau$ for some integer $\tau\geq 0$, and let $\alpha_{1},\alpha_{2},\dots,\alpha_{\tau^{*}}$ denote the $\tau^{*}\leq\tau+1$ distinct transcendental numbers within the set $\left\{\Gamma\left(q\right),\Gamma^{\left(1\right)}\left(q\right),\Gamma^{\left(2\right)}\left(q\right),\ldots,\Gamma^{\left(N\right)}\left(q\right)\right\}$, so that every element of this set lies in $\overline{\mathbb{Q}}\cup\left\{\alpha_{1},\alpha_{2},\dots,\alpha_{\tau^{*}}\right\}$.
+It then follows that every product $\Gamma^{\left(u\right)}\left(q\right)\Gamma^{\left(v\right)}\left(q\right)$ (with $0\leq u,v\leq N$) is contained in the $\overline{\mathbb{Q}}$-vector space
+$\mathcal{U}_{\tau^{*}}=\textrm{span}_{\overline{\mathbb{Q}}}\left\{1,\alpha_{1},\alpha_{2},\dots,\alpha_{\tau^{*}},\alpha_{i}\alpha_{i^{\prime}}:1\leq i\leq i^{\prime}\leq\tau^{*}\right\},$
+implying $c_{j,q}\in\mathcal{U}_{\tau^{*}}$.
+Since the generating set for $\mathcal{U}_{\tau^{*}}$ has cardinality
+$1+\tau^{*}+\frac{\tau^{*}\left(\tau^{*}+1\right)}{2}=\frac{\left(\tau^{*}+1\right)\left(\tau^{*}+2\right)}{2}\leq\frac{\left(\tau+2\right)\left(\tau+3\right)}{2},$
+one can see that
+$\dim_{\overline{\mathbb{Q}}}\left(\mathcal{U}_{\tau^{*}}\right)\leq\frac{\left(\tau+2\right)\left(\tau+3\right)}{2}.$
+However, we also know (from arguments similar to those employed in the proof of Theorem 1) that the elements $\left\{c_{0,q},c_{1,q},\dots,c_{\left\lfloor N/2\right\rfloor,q}\right\}$ are $\overline{\mathbb{Q}}$-linearly independent, forcing the span of this set to have dimension $\left\lfloor N/2\right\rfloor+1$.
+Given that this span is contained in $\mathcal{U}_{\tau^{*}}$, it follows that
+$\left\lfloor N/2\right\rfloor+1\leq\dim_{\overline{\mathbb{Q}}}\left(\mathcal{U}_{\tau^{*}}\right)\leq\frac{\left(\tau+2\right)\left(\tau+3\right)}{2}$
+$\Longrightarrow 2\left\lfloor N/2\right\rfloor+2\leq\left(\tau+2\right)\left(\tau+3\right)$
+$\Longrightarrow\tau\geq\sqrt{2\left\lfloor N/2\right\rfloor+9/4}-5/2.$
+Thus,
+$N\delta_{q}\left(N\right)\geq\max\left\{0,\left\lceil\sqrt{2\left\lfloor N/2\right\rfloor+9/4}-5/2\right\rceil\right\}\geq\max\left\{0,\sqrt{N}-5/2\right\}$
+$\Longrightarrow\delta_{q}\left(N\right)\geq\beta\left(N\right)=\dfrac{\max\left\{0,\sqrt{N}-5/2\right\}}{N},$
+where $\beta\left(N\right)\asymp N^{-1/2}$ as $N\rightarrow\infty$. ∎
+Although the indicated lower bound ($\beta\left(N\right)$) is unaffected by $q$, this does not imply that $\delta_{q}\left(N\right)$ itself is constant over $q\in\frac{1}{2}\mathbb{Z}\setminus\mathbb{Z}_{\leq 0}$.
+After all, $\beta\left(N\right)\asymp N^{-1/2}$ converges to 0 in the limit, thus providing a rather weak (and not very informative) lower bound.
+At the same time, it is worth noting that this asymptotic bound is actually stronger (i.e., larger in magnitude) than the posited asymptotic lower bound of $\widetilde{\beta}\left(N\right)\asymp\ln\left(N\right)/N$ given by Theorem 1(iv) of Powers [10].
+[^6] It also may be of interest to note that $\beta\left(N\right)$ is larger in magnitude than the asymptotic lower bound for the density of irrational odd-$n$ $\zeta\left(n\right)$ among $n\in\left\{3,5,\ldots,N\right\}$ given by Fischler, Sprang, and Zudilin [9]: $\beta_{\zeta}\left(N\right)\asymp\dfrac{2^{\left(1-\varepsilon\right)\ln\left(N\right)/\ln\left(\ln\left(N\right)\right)}}{\left(N-1\right)/2}\asymp N^{\left(1-\varepsilon\right)\ln\left(2\right)/\ln\left(\ln\left(N\right)\right)-1},$ for arbitrarily small $\varepsilon>0$.
+More importantly, the new bound is unconditional, whereas the earlier one is conditional on an unproved (albeit computationally supported) conjecture.
+3. Values of $q\in\mathbb{Q}\setminus\tfrac{1}{2}\mathbb{Z}$
+In this section, we consider transcendental elements in sequences $\left\{\Gamma^{\left(n\right)}\left(q\right)\right\}_{n\geq 1}$ for $q\in\mathbb{Q}\setminus\tfrac{1}{2}\mathbb{Z}$.
+Although the approach is similar to that of Section 2.1, it is restricted by the fact that Euler’s reflection formula (in conjunction with the functional equation) does not yield a simple closed form for $\Gamma\left(q+t\right)\Gamma\left(q-t\right)$ in this case.
+Instead, the reflection formula naturally couples the points $q$ and 1q1-q, leading to a disjunctive result for the sequences $\left\{\Gamma^{\left(n\right)}\left(q\right)\right\}_{n\geq 1}$ and $\left\{\Gamma^{\left(n\right)}\left(1-q\right)\right\}_{n\geq 1}$.
 Theorem 3.
@@
 4. Conclusion
-The present study has pursued a focused objective: generalizing the “infinitely often” transcendence results of Powers [2] to $\Gamma^{\left(n\right)}\left(q\right)$ for values $q\in\mathbb{Q}\setminus\mathbb{Z}_{\leq 0}$ beyond $q=1$.
-This was accomplished for $q\in\tfrac{1}{2}\mathbb{Z}\setminus\mathbb{Z}_{\leq 0}$ in Theorem 1, and partially accomplished for $q\in\mathbb{Q}\setminus\tfrac{1}{2}\mathbb{Z}$ in Theorem 2.
-Natural directions for further research include: (i) improving Theorem 2 by removing the disjunctive limitation; and (ii) ascertaining lower bounds on the asymptotic densities of the transcendental terms $\Gamma^{\left(n\right)}\left(q\right)$ for $q\in\tfrac{1}{2}\mathbb{Z}\setminus\mathbb{Z}_{\leq 0}$.
-The latter objective, although challenging, appears feasible because Theorem 1(iv) of Powers [2] was able to provide such a lower bound for the sequence $\left\{\Gamma^{\left(n\right)}\left(1\right)\right\}_{n\geq 1}$, conditional on a conjecture rigorously certified by computer for all $n\in\left\{2,3,\ldots,26\right\}$.
+In the present study, we investigated the “infinitely often” transcendence of $\Gamma^{\left(n\right)}\left(q\right)$ for $q\in\mathbb{Q}\setminus\mathbb{Z}_{\leq 0}$.
+Theorem 1 showed that for all $q\in\tfrac{1}{2}\mathbb{Z}\setminus\mathbb{Z}_{\leq 0}$, the sequence $\left\{\Gamma^{\left(n\right)}\left(q\right)\right\}_{n\geq 1}$ contains infinitely many transcendental elements, and Theorem 2 gave a lower bound, $\beta\left(N\right)\asymp N^{-1/2}$, for the density of transcendental elements $\Gamma^{\left(n\right)}\left(q\right)$ among $n\in\left\{1,2,\ldots,N\right\}$.
+Although it was not possible to extend the full implications of Theorem 1 to $q\in\mathbb{Q}\setminus\tfrac{1}{2}\mathbb{Z}$, we were able to provide a weaker (disjunctive) result, showing that at least one of the sequences $\left\{\Gamma^{\left(n\right)}\left(q\right)\right\}_{n\geq 1}$, $\left\{\Gamma^{\left(n\right)}\left(1-q\right)\right\}_{n\geq 1}$ contains infinitely many transcendental elements.
+Natural directions for further research include: (i) improving Theorem 3 by removing the disjunctive limitation; and (ii) for fixed $n\in\mathbb{Z}_{\geq 1}$ and $\kappa\in\left(0,1\right)\cap\mathbb{Q}$, ascertaining lower bounds on the densities of transcendental elements $\Gamma^{\left(n\right)}\left(m+\kappa\right)$ among $\left|m\right|\in\left\{0,1,\ldots,M\right\}$.
+Preliminary progress toward the latter objective is presented in Powers [14].

### v3

@@
-“Infinitely Often” Transcendence of Gamma-Function Derivatives
+arXiv:2601.18474v3[math.NT] 27 Feb 2026
 Abstract.
@@
 That result is now generalized to all sequences $\left\{\Gamma^{\left(n\right)}\left(q\right)\right\}_{n\geq 1}$ for $q\in\tfrac{1}{2}\mathbb{Z}\setminus\mathbb{Z}_{\leq 0}$.
-Moreover, for all such $q$ we derive a lower bound, $\beta\left(N\right)=\max\left\{0,\sqrt{N}-5/2\right\}/N$, for the density of transcendental elements $\Gamma^{\left(n\right)}\left(q\right)$ among $n\in\left\{1,2,\ldots,N\right\}$.
+Moreover, for all such $q$ we derive a lower bound, $\beta\left(N\right)=\max\left\{0,\sqrt{N}-5/2\right\}/N$, for the density of transcendental elements $\Gamma^{\left(n\right)}\left(q\right)$ among $n\in\left\{1,2,\ldots,N\right\}$, where $\beta\left(N\right)\asymp N^{-1/2}\rightarrow 0$ as $N\rightarrow\infty$.
 For $q\in\mathbb{Q}\setminus\tfrac{1}{2}\mathbb{Z}$, we find the somewhat weaker result that at least one of the sequences $\left\{\Gamma^{\left(n\right)}\left(q\right)\right\}_{n\geq 1}$, $\left\{\Gamma^{\left(n\right)}\left(1-q\right)\right\}_{n\geq 1}$ contains infinitely many transcendental elements.
@@
 Although it was not possible to extend the full implications of Theorem 1 to $q\in\mathbb{Q}\setminus\tfrac{1}{2}\mathbb{Z}$, we were able to provide a weaker (disjunctive) result, showing that at least one of the sequences $\left\{\Gamma^{\left(n\right)}\left(q\right)\right\}_{n\geq 1}$, $\left\{\Gamma^{\left(n\right)}\left(1-q\right)\right\}_{n\geq 1}$ contains infinitely many transcendental elements.
-Natural directions for further research include: (i) improving Theorem 3 by removing the disjunctive limitation; and (ii) for fixed $n\in\mathbb{Z}_{\geq 1}$ and $\kappa\in\left(0,1\right)\cap\mathbb{Q}$, ascertaining lower bounds on the densities of transcendental elements $\Gamma^{\left(n\right)}\left(m+\kappa\right)$ among $\left|m\right|\in\left\{0,1,\ldots,M\right\}$.
-Preliminary progress toward the latter objective is presented in Powers [14].
+Natural directions for further research include: (i) improving Theorem 3 by removing the disjunctive limitation; and (ii) for fixed $n\in\mathbb{Z}_{\geq 1}$ and $\kappa\in\left[0,1\right)\cap\mathbb{Q}$, ascertaining lower bounds on the densities of transcendental elements $\Gamma^{\left(n\right)}\left(m+\kappa\right)$ among $\left|m\right|\in\left\{0,1,\ldots,M\right\}$.
+Progress toward the latter objective is presented in Powers [14].

### v4

@@
-arXiv:2601.18474v3[math.NT] 27 Feb 2026
+“Infinitely Often” Transcendence of Gamma-Function Derivatives
 Abstract.
@@
 These identities can be iterated to yield the systems [^1] See, e.g., Choi and Srivastava [4].
-$\begin{array}[]{c}\Gamma^{\left(1\right)}\left(1\right)=-\gamma,\\ \Gamma^{\left(2\right)}\left(1\right)=\gamma^{2}+\zeta\left(2\right),\\ \Gamma^{\left(3\right)}\left(1\right)=-\gamma^{3}-3\zeta\left(2\right)\gamma-2\zeta\left(3\right),\\ \Gamma^{\left(4\right)}\left(1\right)=\gamma^{4}+6\zeta\left(2\right)\gamma^{2}+8\zeta\left(3\right)\gamma+\dfrac{27}{2}\zeta\left(4\right),\\ \Gamma^{\left(5\right)}\left(1\right)=-\gamma^{5}-10\zeta\left(2\right)\gamma^{3}-20\zeta\left(3\right)\gamma^{2}-\dfrac{135}{2}\zeta\left(4\right)\gamma-20\zeta\left(2\right)\zeta\left(3\right)-24\zeta\left(5\right),\\$ $\vdots\end{array}$ (2)
+$\begin{array}[]{c}\Gamma^{\left(1\right)}\left(1\right)=-\gamma,\\ \Gamma^{\left(2\right)}\left(1\right)=\gamma^{2}+\zeta\left(2\right),\\ \Gamma^{\left(3\right)}\left(1\right)=-\gamma^{3}-3\zeta\left(2\right)\gamma-2\zeta\left(3\right),\\ \Gamma^{\left(4\right)}\left(1\right)=\gamma^{4}+6\zeta\left(2\right)\gamma^{2}+8\zeta\left(3\right)\gamma+\dfrac{27}{2}\zeta\left(4\right),\\ \Gamma^{\left(5\right)}\left(1\right)=-\gamma^{5}-10\zeta\left(2\right)\gamma^{3}-20\zeta\left(3\right)\gamma^{2}-\dfrac{135}{2}\zeta\left(4\right)\gamma-20\zeta\left(2\right)\zeta\left(3\right)-24\zeta\left(5\right),\\$ $\cdots\end{array}$ (2)
 and
-$\begin{array}[]{c}\Gamma^{\left(1\right)}\left(\tfrac{1}{2}\right)=-\sqrt{\pi}g,\\ \Gamma^{\left(2\right)}\left(\tfrac{1}{2}\right)=\sqrt{\pi}\left(g^{2}+3\zeta\left(2\right)\right),\\ \Gamma^{\left(3\right)}\left(\tfrac{1}{2}\right)=-\sqrt{\pi}\left(g^{3}+9\zeta\left(2\right)g+14\zeta\left(3\right)\right),\\ \Gamma^{\left(4\right)}\left(\tfrac{1}{2}\right)=\sqrt{\pi}\left(g^{4}+18\zeta\left(2\right)g^{2}+56\zeta\left(3\right)g+\dfrac{315}{2}\zeta\left(4\right)\right),\\ \Gamma^{\left(5\right)}\left(\tfrac{1}{2}\right)=-\sqrt{\pi}\left(g^{5}+30\zeta\left(2\right)g^{3}+140\zeta\left(3\right)g^{2}+\dfrac{1575}{2}\zeta\left(4\right)g+420\zeta\left(2\right)\zeta\left(3\right)+744\zeta\left(5\right)\right),\\$ $\vdots\end{array}$ (3)
+$\begin{array}[]{c}\Gamma^{\left(1\right)}\left(\tfrac{1}{2}\right)=-\sqrt{\pi}g,\\ \Gamma^{\left(2\right)}\left(\tfrac{1}{2}\right)=\sqrt{\pi}\left(g^{2}+3\zeta\left(2\right)\right),\\ \Gamma^{\left(3\right)}\left(\tfrac{1}{2}\right)=-\sqrt{\pi}\left(g^{3}+9\zeta\left(2\right)g+14\zeta\left(3\right)\right),\\ \Gamma^{\left(4\right)}\left(\tfrac{1}{2}\right)=\sqrt{\pi}\left(g^{4}+18\zeta\left(2\right)g^{2}+56\zeta\left(3\right)g+\dfrac{315}{2}\zeta\left(4\right)\right),\\ \Gamma^{\left(5\right)}\left(\tfrac{1}{2}\right)=-\sqrt{\pi}\left(g^{5}+30\zeta\left(2\right)g^{3}+140\zeta\left(3\right)g^{2}+\dfrac{1575}{2}\zeta\left(4\right)g+420\zeta\left(2\right)\zeta\left(3\right)+744\zeta\left(5\right)\right),\\ \cdots,\end{array}$ (3)
 where: (i) each $\zeta\left(n\right),\>n\in\left\{2,4,\ldots\right\}$ is a rational multiple of $\pi^{n}$; (ii) $\zeta\left(3\right)$ is known to be irrational (but not necessarily transcendental); [^2] See Apéry [5] and Beukers [6].
@@
 After all, $\beta\left(N\right)\asymp N^{-1/2}$ converges to 0 in the limit, thus providing a rather weak (and not very informative) lower bound.
-At the same time, it is worth noting that this asymptotic bound is actually stronger (i.e., larger in magnitude) than the posited asymptotic lower bound of $\widetilde{\beta}\left(N\right)\asymp\ln\left(N\right)/N$ given by Theorem 1(iv) of Powers [10].
-[^6] It also may be of interest to note that $\beta\left(N\right)$ is larger in magnitude than the asymptotic lower bound for the density of irrational odd-$n$ $\zeta\left(n\right)$ among $n\in\left\{3,5,\ldots,N\right\}$ given by Fischler, Sprang, and Zudilin [9]: $\beta_{\zeta}\left(N\right)\asymp\dfrac{2^{\left(1-\varepsilon\right)\ln\left(N\right)/\ln\left(\ln\left(N\right)\right)}}{\left(N-1\right)/2}\asymp N^{\left(1-\varepsilon\right)\ln\left(2\right)/\ln\left(\ln\left(N\right)\right)-1},$ for arbitrarily small $\varepsilon>0$.
-More importantly, the new bound is unconditional, whereas the earlier one is conditional on an unproved (albeit computationally supported) conjecture.
+[^6] It may be of interest to note that $\beta\left(N\right)$ is larger in magnitude than the asymptotic lower bound for the density of irrational $\zeta\left(n\right)$ among $n\in\left\{3,5,\ldots,N\right\}$ given by Fischler, Sprang, and Zudilin [9]: $\beta_{\zeta}\left(N\right)\asymp\dfrac{2^{\left(1-\varepsilon\right)\ln\left(N\right)/\ln\left(\ln\left(N\right)\right)}}{\left(N-1\right)/2}\asymp N^{\left(1-\varepsilon\right)\ln\left(2\right)/\ln\left(\ln\left(N\right)\right)-1},$ for arbitrarily small $\varepsilon>0$.
 3. Values of $q\in\mathbb{Q}\setminus\tfrac{1}{2}\mathbb{Z}$